# 11affiliationtext: IMT School for Advanced Studies Lucca22affiliationtext: Universidad de Alicante

### v1

@@ Abstract
Inspired by Jang et al. (2022), we propose a Granger causality-in-the-mean test for bivariate $k-$Markov stationary processes based on a recently introduced class of non-linear models, i.e., vine copula models. By means of a simulation study, we show that the proposed test improves on the statistical properties of the original test in Jang et al. (2022), constituting an excellent tool for testing Granger causality in the presence of non-linear dependence structures. Finally, we apply our test to study the pairwise relationships between energy consumption, GDP and investment in the U.S. and, notably, we find that Granger-causality runs two ways between GDP and energy consumption.
@@ 1 Introduction
In this paper, we propose a novel Granger causality-in-the-mean test for bivariate $k-$Markov stationary processes, which is based on vine copula models. Vine copulas are highly flexible models that allow for computationally tractable estimation, model selection procedures, and the representation of a wide array of different joint distributions (Czado and Nagler, 2022). Formally, exploiting the concept of pair copula construction (Joe, 1996), a multivariate copula is constructed using bivariate copulas as building blocks, overcoming the limited and restrictive selection of multivariate copulas. As shown in recent literature (Brechmann and Czado, 2015; Beare and Seo, 2015; Smith, 2015; Nagler et al., 2022), vine copulas are non-linear models that can represent multivariate time series by capturing both their cross-sectional and serial dependence within the same model, giving place to a myriad of applications in the context of time series modelling, one of them being the study of Granger causality.
@@
In this contribution, we start from the test procedure proposed in Jang et al. (2022) and incorporate some relevant methodological differences. Thus, we build our non-linear Granger causality-in-the-mean test for bivariate $k-$Markov stationary processes, which we call M-vine Granger causality test. By means of a simulation study based on the setting from Song and Taamouti (2018), we find that the M-vine test has a higher power than the original version from Jang et al. (2022) in all scenarios considered, whil still controlling its size under or considerably close to the predefined significance level. Moreover, for linear specifications and large samples, our test displays a power that resembles the one from a traditional Granger-causality test based on linear models, even if the latter is based on the real data generating process. Yet, for non-linear specifications, the M-vine test has a substantially higher power than the linear Granger-causality test. Hence, we show that our test improves on the statistical properties of the original one in Jang et al. (2022) and makes it an excellent tool for testing Granger causality in the presence of non-linear dependence structures.
@@
The rest of the paper is organized as follows. Section 2 introduces the methodology and the setting of the M-vine test, while section 3 studies the statistical properties of the proposed test by means of a simulation study. In section 4, we present a macroeconomic application to pairwise relationships energy consumption, GDP and investment in the U.S. Conclusions and final remarks are offered in Section 5.
@@ 2 Methodology
Measures of Granger causality are classified into three types, according to the criterion chosen to measure their prediction quality: namely, Granger causality in the mean, in the quantiles or in the distributions. Here, we focus only on the first, which is the most widely used in literature, while we refer to Lee and Yang (2014) and Candelon and Tokpavi (2016) for the others. Nonetheless, we believe it is important to remark that the methodology that we present can be naturally extended to Granger causality in quantiles (Song and Taamouti, 2021; Hyuna Jang and Noh, 2023).
@@
Definition 1
(Granger causality in the mean) We say that $Y$ causes $X$ in the sense of Granger, if
@@
Jang et al. (2022) proposed a semi-parametric Granger causality test based on the above measure $GC^{mean}(Y\rightarrow X)$, estimated using a copula approach, which is able to capture non-linear dependence while having desirable statistical properties when compared to other existing tests. This test is our starting point. We here propose a test of Granger causality in the mean based on the M-vine copula structure presented in Beare and Seo (2015) for a bivariate $k-$Markov stationary stochastic process. The presented test improves both on the size and the power when compared to the one in Jang et al. (2022), and it is more suitable for empirical applications since it takes into account possible non-linear dependencies that are commonly encountered when working with real data. Please, refer to Appendix A.1 for some recalls on the notion of copulas and, in particular, on vine copulas.
@@ 2.1 M-vine Granger causality test
Let $(X,Y)=(X_{t},Y_{t})_{t}$ be a bivariate $k-$Markov stationary stochastic process and let $\{(x_{t},y_{t})\;:\;t=1,\dots,T\}$ be a sample of it. To simplify notation, from now on, we will assume $k=1$, but everything can be naturally extended to any $k\in\mathbb{N}\setminus\{0\}$.
We can split the proposed test, which we call the M-vine Granger causality test, into two parts, say part A and part B: the first regards the estimation of the Granger causality (in the mean) measure, and part B concerns the computation of the $p$-value of the test.
We start by describing part A. We use the sample version of the above-recalled measure $GC^{mean}(Y\rightarrow X)$, that is
@@
where the quantities $\widehat{\mathbb{E}}[X_{t}|X_{t-1}=x_{t-1}]$ and $\widehat{\mathbb{E}}[X_{t}|X_{t-1}=x_{t-1},Y_{t-1}=y_{t-1}]$ are computed as follows:
@@
    we fit an M-vine copula model and estimate its respective parameters for the observations from the series $X$, that is for $M_{x}=(x_{t})_{t=1\dots,T}$;
@@
    using the model of Step 1), for each $t=T_{0},\dots,T$, we generate $N$ i.i.d. predictions $\{\tilde{x}^{M_{x}}_{i,t}\}_{i=1}^{N}$ of $X_{t}$ given $X_{t-1}=x_{t-1}$, so that we obtain
@@
    repeat Step 1) and Step 2) for the observations from the bivariate series $(X,Y)$, that is for $M_{x,y}=(x_{t},y_{t})_{t=1\dots,T}$, in order to obtain
    $\widehat{\mathbb{E}}[X_{t}|X_{t-1}=x_{t-1},Y_{t-1}=y_{t-1}]=\frac{1}{N}\sum_{i=1}^{N}\tilde{x}^{M_{x,y}}_{i,t}\qquad t=T_{0},\dots,T.$
@@
Thus, roughly speaking, a sample estimate of Granger causality in the mean, i.e. $\widehat{GC}^{mean}_{Y\rightarrow X}$, measures the log-difference between the prediction error related to the model with an M-vine copula structure fitted only on the information set from $X$, and the prediction error computed with the M-vine copula model fitted to the entire sample form both $X$ and $Y$, that is using the information set of both series. In line with the definition of Granger causality in the mean, if this estimated quantity is significantly higher than zero, we reject the null hypothesis of no Granger causality (in the mean) from $Y$ to $X$: indeed, under the null hypothesis, the measure $GC^{mean}(Y\rightarrow X)$ is zero, and so, the higher is the value of the statistics $\widehat{GC}^{mean}_{Y\rightarrow X}$, the more statistically significant is the evidence of the presence of Granger causality (in the mean) running from $Y$ to $X$. Regarding $T_{0}\geq k+1$, we can choose it by taking into account the goodness of fit of the models to the data.
In order to test if $\widehat{GC}^{mean}_{Y\rightarrow X}$ is statistically greater than zero, we rely, as in Jang et al. (2022), on a bootstrap method, which takes inspiration from Paparoditis and Politis (2000). The procedure relies on the M-vine copula model fitted on the entire sample $M_{x,y}=\{(x_{t},y_{t}):t=1,\dots,T\}$ in order to generate bootstrapped samples under the null hypothesis and use them for computing the $p-$value for the test. More precisely, the part B of the proposed procedure works as follows:
@@
    from the first tree of the found M-vine copula structure, extract, for each $t=1,\dots,T$, both the copula between $X_{t}$ and $X_{t+1}$, say $c_{X_{t},X_{t+1}}$, related to the conditional distribution of $X_{t+1}$ given $X_{t}$, and the copula between $X_{t}$ and $Y_{t}$, say $c_{X_{t},Y_{t}}$, related to the conditional distribution of $Y_{t}$ given $X_{t}$;
@@
    using this bootstrapped sample $\{(x_{t}^{0},y_{t}^{0}):t=1,\dots,T\}$, compute the bootstrapped version of the test statistic under the null hypothesis, say $\widehat{GC}^{0\;mean}_{Y\rightarrow X}$;
@@
    repeat the above steps $B$ times, so that we get $B$ bootstrapped samples and so $\widehat{GC}^{0\;mean}_{j\;Y\rightarrow X}$ for $j=1,\dots,B$;
@@
    compute the bootstrapped $p$-value for the test by
@@
There are two relevant differences between the test by Jang et al. (2022) and the one we propose. Firstly, in the above-described Step 1A), we choose the class of the M-vine copulas to fit the data. The same model is then used in the bootstrap procedure to construct the $p$-value for the test. On the contrary, Jang et al. (2022) use a more general class of copula models, i.e., the S-vine copulas, but then they restrict to the M-vine copulas in the bootstrap procedure (part B). The justification of this fact given in Jang et al. (2022) simply relies on the implementation of the bootstrap procedure: indeed, to draw the samples in the above Steps B2) and B3), both $c_{X_{t},X_{t+1}}$ and $c_{X_{t},Y_{t}}$ are needed, and these two copulas are guaranteed to be present in the first tree of a given M-vine structure, whilst these two copulas might not appear in an S-vine structure. Consequently, in Jang et al. (2022), the conditional distribution used for computing the bootstrapped sample of the Granger causality might be different from the ones in the model fitted to the data to estimate the Granger causality measure. In our procedure, the two parts, A and B, of the Granger causality test are aligned: they rely on the same class of vine copulas.
Secondly, in the above Step 1A), we fit the M-vine copula model on the entire sample, i.e., on all the $T$ observations, while Jang et al. (2022) split the entire sample into a training set, say observations from $t=1$ to $t=T^{*}$ (usually $T^{*}=T/2$), and a testing set, say observations from $t=T^{*}+1$ to $t=T$. Therefore, the fit of the $S$-vine copula model is performed only on the training set, while the testing set is used to compute the Granger causality statistics. The latter choice implies that all the predictions $\{\tilde{x}^{M_{x}}_{i,t}\}_{i=1}^{N}$ and $\tilde{x}^{M_{x,y}}_{i,t}$ for $t=T^{*}+1,\dots,t$, used to compute the Granger causality statistics are based on a model fitted only on the observations from $t=1,\dots,T^{*}$. In real applications, when the time series are not perfectly stationary or the sample is small, this way of proceeding could generate poor estimates of the Granger causality measure. This is especially true in the case of non-linearities in the data generating process. In our case, instead, predictions are based on a model fitted to the entire sample, and this potentially makes our test better at detecting Granger causality, especially when the dependence structure is not linear. In the next subsection, we show that our variant has a better performance in terms of size and power.
@@ 3 Simulation Study
In this section, we analyse the statistical performance of our M-vine Granger causality test (briefly, M-vine test) and compare it, in terms of size and power, with the traditional linear Granger-causality test based on VAR models  [^1] The linear test refers to a Wald test based on restricted and unrestricted VAR models, and their results are obtained after application of grangertest from the R package lmtest. and the S-vine Granger causality test (S-vine test) by Jang et al. (2022). We recall that, as shown in Jang et al. (2022), the last one has a good statistical performance in terms of size and power when compared to other non-linear Granger causality tests (Diks and Panchenko, 2006; Song and Taamouti, 2018; Kim et al., 2020). Specifically, through a simulation study, they show that S-vine test, ST test (Song and Taamouti, 2018) and KLH test (Kim et al., 2020) are the ones whose size is closer to the predefined significance level, whereas the size of DP test (Diks and Panchenko, 2006) is considerably smaller than the significance level and this suggests a tendency of this test to under-reject. Indeed, in terms of statistical power, Jang et al. (2022) evidences that DP test is the one that performs the worst. On the contrary, S-vine test shows, in large samples ($T\geq 200$ for Markov processes of order $k=1$), a good power for all the dependence structure, while KLH test has an excellent power only when working with linear or quasi-linear models and ST test displays a good power exclusively when the dependence structure is non-linear. In the following, we will show that our M-vine test outperforms S-vine test in all the considered scenarios, exhibiting a very good power already starting from $T=100$ in the case of Markor order $k=1$.
We performed a simulation study based on the assessment models from Song and Taamouti (2018). We used their three size assessment models and their three power assessment models while incorporating an additional fourth power assessment model where both series come from non-linear specifications:
@@
For the simulation study, we used $S=500$ simulations for each model with sample sizes of $T=50$, $T=100$ and $T=200$, and computed the empirical size and power using a significance level $\alpha=0.05$. Furthermore, for part A of the procedure, we used $N=200$ i.i.d. predictions for each series to estimate the conditional expectations and, for part B, i.e. the bootstrapping procedure to obtain the $p$-values, we used a number $B=200$ of bootstrapped samples. We used $T_{0}=T/2$ in the computation of the test statistics in accordance with the choice $T^{*}=T/2$ in Jang et al. (2022).
@@
S1 0.046 0.066 0.058 0.050 0.070 0.048 0.056 0.050 0.058
S2 0.044 0.052 0.058 0.052 0.080 0.054 0.062 0.066 0.056
S3 0.058 0.070 0.060 0.054 0.054 0.052 0.050 0.070 0.046

### v2

@@ Abstract
Inspired by Jang et al. (2022), we propose a Granger causality-in-the-mean test for bivariate $k-$Markov stationary processes based on a recently introduced class of non-linear models, i.e., vine copula models. By means of a simulation study, we show that the proposed test improves on the statistical properties of the original test in Jang et al. (2022), and also of other previous methods, constituting an excellent tool for testing Granger causality in the presence of non-linear dependence structures. Finally, we apply our test to study the pairwise relationships between energy consumption, GDP and investment in the U.S. and, notably, we find that Granger-causality runs two ways between GDP and energy consumption.
@@ 1 Introduction
In this paper, we propose a novel Granger causality-in-the-mean test for bivariate $k-$Markov stationary processes, which is based on vine copula models. In general, copula functions are the mathematical tool that describes any type of dependence structure and so they are able to capture non-linear dependencies. In particular, vine copulas are highly flexible models that allow for the representation of a wide array of different joint distributions (Czado and Nagler, 2022) and computationally tractable model-selection and estimation procedures. Formally, exploiting the concept of pair copula construction (Joe, 1996), a multivariate copula is constructed using bivariate copulas as building blocks, overcoming the limited and restrictive selection of multivariate copulas. As shown in recent literature (Brechmann and Czado, 2015; Beare and Seo, 2015; Smith, 2015; Nagler et al., 2022), vine copulas are non-linear models that can represent multivariate time series by capturing both their cross-sectional and serial dependence within the same model, giving place to a myriad of applications in the context of time series modelling, one of them being the study of Granger causality.
@@
In this contribution, we start from the test procedure proposed in Jang et al. (2022) and incorporate some relevant methodological differences. Thus, we build our non-linear Granger causality-in-the-mean test for bivariate $k-$Markov stationary processes, which we call M-vine Granger causality test. By means of a simulation study based on the setting from Song and Taamouti (2018), we find that the M-vine test has a higher power than the original version from Jang et al. (2022) in all scenarios considered, while still controlling its size under or considerably close to the predefined significance level. Moreover, for linear specifications and large samples, our test displays a power that resembles the one from a traditional Granger-causality test based on linear models, even if the latter is based on the real data generating process. Yet, for non-linear specifications, the M-vine test has a substantially higher power than the linear Granger-causality test. Hence, we show that our test improves on the statistical properties of the original one in Jang et al. (2022), and also of other previous methods, and makes it an excellent tool for testing Granger causality in the presence of non-linear dependence structures.
@@
The rest of the paper is organized as follows. Section 2 introduces the methodology and the setting of the M-vine test, while section 3 studies the statistical properties of the proposed test by means of a simulation study. In section 4, we present a macroeconomic application to analyse the pairwise relationships between energy consumption, GDP and investment in the U.S. Conclusions and final remarks are offered in Section 5.
@@ 2 Methodology
Measures of Granger causality are classified into three types, according to the criterion chosen to measure their prediction quality: namely, Granger causality in the mean, in the quantiles or in distribution. Here, we focus only on the first, which is the most widely used in literature, while we refer to Lee and Yang (2014) and Candelon and Tokpavi (2016) for the others. Nonetheless, we believe it is important to remark that the methodology that we present can be naturally extended to Granger causality in quantiles (Song and Taamouti, 2021; Hyuna Jang and Noh, 2023).
@@
Definition 1.
(Granger causality in the mean) Assuming $X_{t}$ and $Y_{t}$ to be square integrable, we say that $Y$ causes $X$ in the sense of Granger, if
@@
Jang et al. (2022) proposed a semi-parametric Granger causality test based on the above measure $GC^{mean}(Y\rightarrow X)$, estimated using a copula approach (see Appendix A.1). Copula functions are the mathematical tool that describes any type of dependence structure. Therefore, this test is able to capture non-linear dependence and, moreover, it exhibits good statistical properties when compared to other existing non-linear tests (see Jang et al. (2022)). Hence, it is our starting point. We here propose a test for Granger causality in the mean based on the M-vine copula structure presented in Beare and Seo (2015) for a bivariate $k-$Markov stationary stochastic process. The presented test improves both on the size and the power when compared to the one in Jang et al. (2022), and it is more suitable for empirical applications since it takes into account possible non-linear dependencies that are commonly encountered when working with real data. Please, refer to Appendix A.1 for some recalls on the notion of copulas and, in particular, on vine copulas.
@@ 2.1 M-vine Granger causality test
Let $(X,Y)=(X_{t},Y_{t})_{t}$ be a bivariate $k-$Markov square-integrable stationary stochastic process and let $\{(x_{t},y_{t})\;:\;t=1,\dots,T\}$ be a sample of it. To simplify notation, from now on, we will assume $k=1$, but everything can be naturally extended to any $k\in\mathbb{N}\setminus\{0\}$.
We can split the proposed test, which we call the M-vine Granger causality test, into two parts, say part A and part B: the first one regards the evaluation of an estimator of the Granger causality (in the mean) measure, which is used as the test statistics, and the second one concerns the computation of the $p$-value of the test.
We start by describing part A (computation of the value of the test statistic). We use the sample version of the above-recalled measure $GC^{mean}(Y\rightarrow X)$, that is
@@
where the quantities $\widehat{\mathbb{E}}[X_{t}|X_{t-1}=x_{t-1}]$ and $\widehat{\mathbb{E}}[X_{t}|X_{t-1}=x_{t-1},Y_{t-1}=y_{t-1}]$ are computed by fitting an M-vine copula model to the data (see Appendix A.1.2) and using it to generate i.i.d. observations of $X_{t}$ given $X_{t-1}=x_{t-1}$ and of $X_{t}$ given $X_{t-1}=x_{t-1}$ and $Y_{t-1}=y_{t-1}$ and using the corresponding empirical means to estimate the needed conditional expectations. More precisely, we proceed as follows:
@@
    we fit an M-vine copula model and estimate its respective parameters for the observations from the series $X$, that is for $M_{x}=(x_{t})_{t=1\dots,T}$ we adopt the selection and estimation procedure introduced and studied in Nagler et al. (2022), where, among other things, its consistency is proven under suitable assumptions; see Section A.1.3 for details);
@@
    using the model obtained from Step A1), for each $t=T_{0},\dots,T$, we generate $N$ i.i.d. predictions $\{\tilde{x}^{M_{x}}_{i,t}\}_{i=1}^{N}$ of $X_{t}$ given $X_{t-1}=x_{t-1}$, so that we can estimate the conditional mean of $X_{t}$ given $X_{t-1}=x_{t-1}$ by the empirical mean of the sample $\{\tilde{x}^{M_{x}}_{i,t}\}_{i=1}^{N}$, that is we set
@@
    repeat Step A1) and Step A2) for the observations from the bivariate series $(X,Y)$, that is for $M_{xy}=(x_{t},y_{t})_{t=1\dots,T}$, in order to obtain
    $\widehat{\mathbb{E}}[X_{t}|X_{t-1}=x_{t-1},Y_{t-1}=y_{t-1}]=\frac{1}{N}\sum_{i=1}^{N}\tilde{x}^{M_{xy}}_{i,t}\qquad t=T_{0},\dots,T.$
@@
Note that, by the strong law of large numbers, for $N\to+\infty$, the empirical means $\frac{1}{N}\sum_{i=1}^{N}\tilde{x}^{M_{x}}_{i,t}$ and $\frac{1}{N}\sum_{i=1}^{N}\tilde{x}^{M_{xy}}_{i,t}$ are strongly consistent estimators of the conditional mean of $X_{t}$ given $X_{t-1}=x_{t-1}$ and given $X_{t-1}=x_{t-1},\,Y_{t-1}=y_{t-1}$ respectively, with respect to the model obtained from Step A1. Therefore, the computed quantity $\widehat{GC}^{mean}_{Y\rightarrow X}$ results to be a strongly consistent estimator (for $N\to+\infty$) of the Granger causality measure with respect to the model obtained from Step A1 applied to $M_{x}$ and $M_{xy}$).
Roughly speaking, $\widehat{GC}^{mean}_{Y\rightarrow X}$, measures the log-difference between the prediction error related to the model with an M-vine copula structure fitted only on the information set from $X$, and the prediction error computed with the M-vine copula model fitted to the entire sample form both $X$ and $Y$, that is using the information set of both series. In line with the definition of Granger causality in the mean, if this estimated quantity is significantly higher than zero, we reject the null hypothesis of no Granger causality (in the mean) from $Y$ to $X$: indeed, under the null hypothesis, the measure $GC^{mean}(Y\rightarrow X)$ is zero, and so, the higher is the value of the statistics $\widehat{GC}^{mean}_{Y\rightarrow X}$, the more statistically significant is the evidence of the presence of Granger causality (in the mean) running from $Y$ to $X$. Regarding $T_{0}\geq k+1$, we can choose it by taking into account the goodness of fit of the models to the data.
In order to test if $\widehat{GC}^{mean}_{Y\rightarrow X}$ is statistically greater than zero, we rely, as in Jang et al. (2022), on a method, which takes inspiration from Paparoditis and Politis (2000). This method relies on the M-vine copula model fitted on the entire sample $M_{xy}=\{(x_{t},y_{t}):t=1,\dots,T\}$ in order to generate independent samples under the null hypothesis of no Granger causality (in the mean) and use them for computing the $p-$value for the test. More precisely, the part B (computation of the $p$-value) of the proposed procedure works as follows:
@@
    from the first tree of the obtained M-vine copula structure, extract, for each $t=1,\dots,T$, both the copula between $X_{t}$ and $X_{t+1}$, say $c_{X_{t},X_{t+1}}$, related to the conditional distribution of $X_{t+1}$ given $X_{t}$, and the copula between $X_{t}$ and $Y_{t}$, say $c_{X_{t},Y_{t}}$, related to the conditional distribution of $Y_{t}$ given $X_{t}$;
@@
    using this generated sample $\{(x_{t}^{0},y_{t}^{0}):t=1,\dots,T\}$, compute the quantity $\widehat{GC}^{0\;mean}_{Y\rightarrow X}$, that gives a simulated value of the test statistic under the null hypothesis;
@@
    repeat the above steps $B$ times, so that we get $B$ independent simulated values under the null hypothesis: $\widehat{GC}^{0\;mean}_{j\;Y\rightarrow X}$ for $j=1,\dots,B$;
@@
    compute the $p$-value for the test by the empirical mean
@@
There are two relevant differences between the test by Jang et al. (2022) and the one we propose. Firstly, in the above-described Step 1A), we choose the class of the M-vine copulas to fit the data and to compute the value of the test statistic. The same model is then used in the procedure (part B) to compute the $p$-value for the test. On the contrary, Jang et al. (2022) use a more general class of copula models (that includes the M-vine copulas), i.e. the class of S-vine copulas, for fitting the data and computing the test statistic, but then they restrict to the M-vine copulas for the computation of the $p$-value. The justification of this fact given in Jang et al. (2022) simply relies on the implementation of this last computation: indeed, to draw the samples in the above Steps B2) and B3), both $c_{X_{t},X_{t+1}}$ and $c_{X_{t},Y_{t}}$ are needed, and these two copulas are guaranteed to be present in the first tree of a given M-vine structure, whilst these two copulas might not appear in an S-vine structure (see Section A.1.2 for more details). Consequently, in Jang et al. (2022), the conditional distributions used for generating a sample of the Granger causality measure under the null hypothesis might be different from the ones in the model fitted to the data to estimate the Granger causality measure, i.e. to compute the value of the test statistic. In our procedure, the two parts, A (computation of the test statistic) and B (computation of the $p$-value), of the Granger causality test are aligned: they rely on the same class of vine copulas, i.e. the M-vine copulas.
Secondly, in the above Step 1A), we fit the M-vine copula model on the entire sample, i.e., on all the $T$ observations, while Jang et al. (2022) split the entire sample into a training set, say observations from $t=1$ to $t=T^{*}$ (usually $T^{*}=T/2$), and a testing set, say observations from $t=T^{*}+1$ to $t=T$. Therefore, the fit of the $S$-vine copula model is performed only on the training set, while the testing set is used to compute the Granger causality statistics. The latter choice implies that all the predictions $\{\tilde{x}^{M_{x}}_{i,t}\}_{i=1}^{N}$ and $\tilde{x}^{M_{xy}}_{i,t}$ for $t=T^{*}+1,\dots,t$, used to compute the Granger causality statistics are based on a model fitted only on the observations from $t=1,\dots,T^{*}$. In real applications, when the time series are not perfectly stationary or the sample is small, this way of proceeding could generate poor estimates of the Granger causality measure. This is especially true in the case of non-linearities in the data generating process. In our case, instead, predictions are based on a model fitted to the entire sample, and this potentially makes our test better at detecting Granger causality, especially when the dependence structure is not linear. Note that fitting the model on more observations does not mean to add more model parameters, hence the risk of “overfitting” the data is not present.
In the next subsection, we show that our variant has a better performance in terms of size and power.
@@ 3 Simulation Study
In this section, we analyse the statistical performance of our M-vine Granger causality test (briefly, M-vine test) and compare it, in terms of size and power, with the traditional linear Granger-causality test based on VAR models  [^1] The results of the linear test are obtained using the function grangertest from the R package lmtest. For the formulation of the linear test, we refer to Appendix A.2, the S-vine Granger causality test (briefly, S-vine test) by Jang et al. (2022), and with a non-linear extension of the traditional Granger causality test using feed-forward neural networks from Hmamouche (2020) (briefly, NlinTS)  [^2] For the simulation study, we decided to compare the M-vine test only with methods for which the code is available, so that we could run the simulations ourselves and ensure the validity of the results.. We recall that, as shown in Jang et al. (2022), the S-vine test has a good statistical performance in terms of size and power when compared to other non-linear Granger causality tests (Diks and Panchenko, 2006; Song and Taamouti, 2018; Kim et al., 2020). Specifically, through a simulation study, they show that S-vine test, ST test (Song and Taamouti, 2018) and KLH test (Kim et al., 2020) are the ones whose size is closer to the predefined significance level, whereas the size of the DP test (Diks and Panchenko, 2006) is considerably smaller than the significance level and this suggests a tendency of this test to under-reject. Indeed, in terms of statistical power, Jang et al. (2022) evidences that the DP test is the one that performs the worst. On the contrary, the S-vine test shows, in large samples ($T\geq 200$ for Markov processes of order $k=1$), a good power for all the dependence structure, while the KLH test has an excellent power only when working with linear or quasi-linear models and the ST test displays a good power exclusively when the dependence structure is non-linear. In the following, we will show in particular that our M-vine test outperforms the S-vine test in all the considered scenarios, exhibiting a very good power already starting from $T=100$ in the case of Markov order $k=1$.
We performed a simulation study based on the assessment models from Jang et al. (2022). We used their three size assessment models (S1, S2 and S3) and their three power assessment models (P1, P2 and P3). Furthermore, we enhanced the exercise by including other data generating processes: two size assessment models (S4 and S5) and one power assessment model (P5) taken from Song and Taamouti (2018) and an additional power assessment model (P4) where both series come from non-linear specifications.
@@
For the simulation study, we used $S=500$ simulations for each model with sample sizes of $T=50$, $T=100$ and $T=200$, and computed the empirical size and power using a significance level $\alpha=0.05$. Furthermore, for part A of the procedure, we used $N=200$ i.i.d. predictions for each series to estimate the conditional expectations and, for part B, i.e. the computation of the $p$-value, we used a number $B=200$ of generated samples. We used $T_{0}=T/2$ in the computation of the test statistic in accordance with the choice $T^{*}=T/2$ in Jang et al. (2022).
@@
row1 = c, column4 = c, column8 = c, column12 = c, cell11 = r=2, cell12 = c=4, cell16 = c=4, cell110 = c=4, cell32 = c, cell33 = c, cell35 = c, cell36 = c, cell37 = c, cell39 = c, cell310 = c, cell311 = c, cell313 = c, cell42 = c, cell43 = c, cell45 = c, cell46 = c, cell47 = c, cell49 = c, cell410 = c, cell411 = c, cell413 = c, cell52 = c, cell53 = c, cell55 = c, cell56 = c, cell57 = c, cell59 = c, cell510 = c, cell511 = c, cell513 = c, cell62 = c, cell63 = c, cell65 = c, cell66 = c, cell67 = c, cell69 = c, cell610 = c, cell611 = c, cell613 = c, cell72 = c, cell73 = c, cell75 = c, cell76 = c, cell77 = c, cell79 = c, cell710 = c, cell711 = c, cell713 = c, vline6,10 = 2-7, hline1,8 = -, hline2-3 = 2-13, Model & T=50 T=100 T=200
M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear M-vine Jang et al. NlinTS Linear
S1 0.046 0.066 0.000 0.058 0.050 0.070 0.016 0.048 0.056 0.050 0.106 0.058
S2 0.044 0.052 0.000 0.058 0.052 0.080 0.026 0.054 0.062 0.066 0.074 0.056
S3 0.058 0.070 0.000 0.060 0.054 0.054 0.014 0.052 0.050 0.070 0.090 0.046
S4 0.060 0.082 0.002 0.056 0.060 0.044 0.026 0.064 0.054 0.048 0.080 0.056
S5 0.056 0.074 0.000 0.040 0.054 0.060 0.026 0.066 0.050 0.054 0.072 0.036